Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-h10. What is the labor efficiency variance for March? (Indicate the effect of each variance by selecting F for favorable, U
Answer #1

10) Labor efficiency variance = (30000*2-55000)*14 = 70000 F

11) Labor spending variance = (30000*28)-(55000*15) = 15000 F

12) variable overhead in planning budget = 25000*10 = $250000

13) variable overhead in Flexible budget = 30000*10 = $300000

14) Variable overhead rate variance = (5*55000-280500) = 5500 U

15) Variable overhead efficiency variance = (60000-55000)*5 = 25000 F